# A New Construction Structure on Multi-access Coded Caching with Linear Subpacketization: Cyclic Multi-Access Non-Half-Sum Disjoint Packing

### v1

@@ V-A3 Comparison with the WCWL scheme in [25]
Let $K=q^{n}L$ and $t=q^{n}-\lfloor{q-1}\rfloor^{n}$ in [25]. We compare the WCWL scheme with the proposed scheme in the following cases:
$g_{\text{WCWL}}$ $=2\lfloor\frac{K}{K-tL+1}\rfloor=2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor.$
@@
$\frac{g_{\text{WCWL}}}{g}$ $=\frac{2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor}{2^{n}}=2^{1-n}\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor$
@@
    $F_{\text{WCWL}}=K=q^{n}L,R_{\text{WCWL}}=\frac{(K-tL)(K-tL+1)}{2K}=\frac{\lfloor{q-1}\rfloor^{n}\left(L\lfloor{q-1}\rfloor^{n}+1\right)}{2q^{n}}.$
@@
    $\frac{F_{\text{WCWL}}}{F}=\frac{q^{n}L}{q^{n}L}=1,\quad\frac{R_{\text{WCWL}}}{R}=\frac{\frac{\lfloor{q-1}\rfloor^{n}\left(L\lfloor{q-1}\rfloor^{n}+1\right)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=2^{n}\frac{L\lfloor{q-1}\rfloor^{n}+1}{2q^{n}L}\textgreater\frac{2^{n}(q-1)^{n}}{2q^{n}}=\frac{1}{2}\cdot(2-\frac{2}{q})^{n}.$
@@
    $F_{\text{WCWL}}=\left(2\lfloor\frac{K}{K-tL+1}\rfloor+1\right)K,\ \ \ R_{\text{WCWL}}=\frac{K-tL}{2\lfloor\frac{K}{K-t+1}\rfloor+1}=\frac{\lfloor{q-1}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor+1}.$
@@
    $\frac{R_{\text{WCWL}}}{R}$ $=\frac{\frac{\lfloor{q-1}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor+1}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}}{2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor+1}\geq\frac{2^{n}(q-1)^{n}L}{2\left(2q^{n}+(q-1)^{n}\right)L}=\frac{2^{n}}{2(\frac{q}{{q-1}})^{n}+1}.$
  • Otherwise, the WCWL scheme can be obtain as follow.
    $F_{\text{WCWL}}=2\left\lfloor\frac{K}{K-tL+1}\right\rfloor K,\ \ \ R_{\text{WCWL}}=\frac{K-tL}{2\lfloor\frac{K}{K-tL+1}\rfloor}={\frac{\lfloor{q-1}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor}}.$
@@
    $\frac{F_{\text{WCWL}}}{F}=\frac{2\left\lfloor\frac{K}{K-tL+1}\right\rfloor K}{q^{n}L}=2\left\lfloor\frac{K}{K-tL+1}\right\rfloor,\quad\frac{R_{\text{WCWL}}}{R}=\frac{\frac{\lfloor{q-1}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}}{2\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor}>\frac{2^{n}}{2(\frac{q}{{q-1}})^{n}}=\frac{1}{2}\cdot(2-\frac{2}{q})^{n}.$
As a result, compared to the WCWL scheme in [25], our scheme obtaining the same subpacketization or the reduction amount of our subpacketization is larger than $2$ times. The multiplicative reduction amount by our transmission is at least $\frac{2^{n}}{2(\frac{q}{{q-1}})^{n}+1}$. When $K=9$ or $K\geq 3^{3}L$, the formula $\frac{2^{n}}{2(\frac{q}{{q-1}})^{n}+1}\geq 1$ always holds. The coded caching gain of our scheme is $2^{1-n}\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor$ times that of the WCWL scheme. Since $K\geq 3^{n}L$, $2^{1-n}\left\lfloor\frac{q^{n}L}{\lfloor{q-1}\rfloor^{n}L+1}\right\rfloor$ is always less than or equal to 1.
@@ V-A4 Comparison with the SR2 scheme in [28]
Let $K=q^{n}L$ and $k=q^{n}-\lfloor{q-1}\rfloor^{n}$ in [28], i.e., the $10$th row of Table I. We compare the SR2 scheme with the proposed scheme in the following cases:
$F_{\text{SR2}}=K=q^{n}L,\quad R_{\text{SR2}}=\frac{(K-kL)(K-kL+k)}{2K}=\frac{\lfloor{q-1}\rfloor^{n}\left(\lfloor{q-1}\rfloor^{n}L+q^{n}-\lfloor{q-1}\rfloor^{n}\right)}{2q^{n}}.$
@@
$\frac{F_{\text{SR2}}}{F}=\frac{q^{n}L}{q^{n}L}=1,\quad\frac{R_{\text{SR2}}}{R}=\frac{\frac{\lfloor{q-1}\rfloor^{n}\left(\lfloor{q-1}\rfloor^{n}L+q^{n}-\lfloor{q-1}\rfloor^{n}\right)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}\lfloor{q-1}\rfloor^{n}(L-1)+q^{n}}{2q^{n}L}.$
As a result, compared to the SR2 scheme in [28], our scheme either obtaining the same subpacketization. The multiplicative reduction amount by our transmission is at least $\frac{2^{n}\lfloor{q-1}\rfloor^{n}(L-1)+q^{n}}{2q^{n}L}$. When $K\geq 3q^{2}$ ,the formula $\frac{2^{n}\lfloor{q-1}\rfloor^{n}(L-1)+q^{n}}{2q^{n}L}\geq 1$ always holds.
@@ V-B Numerical comparisons
V-B1 Comparison with the schemes in [8, 22, 25, 24]
When $K=839$ and $L=3$, we have our scheme and the schemes in [8, 22, 23, 24] are shown in Fig. 3(a) and Fig. 3(b). It can be seen that our scheme can achieve a lower rate while maintaining a lower or same subpacketization than the schemes in [22, 25, 24], and achieve a lower subpacketization and a slightly higher rate than the schemes in [8].
(a) Memory ratio-load tradeoff for $K=839$,$L=3$
(b) Memory ratio-subpacketization tradeoff for $K=839$,$L=3$
@@ V-B1 Comparison with the schemes in [8, 22, 25, 24]
V-B2 Comparison with the schemes in [29, 26, 28, 27, 30]
The WCWL scheme is superior to these schemes. Furthrtmore, the subpacketization of our scheme is always equal to or smaller than WCWL scheme, and the coded caching gain is always greater than or equal to that of WCWL scheme, so our scheme is also outperforms these schemes.
@@ VI Conclusion
This paper proposes a novel combinational structure called as cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP), which is an extension of the traditional NHSDP structure to MACC system. The CMA-NHSDP can realize a multi-access coded caching scheme with linear subpacketization while minimizing transmission load. In terms of performance comparison, compared to some existing schemes with linear subpacketization, our scheme has larger coded caching gain, thus achieving a lower rate. And our scheme achieves far lower subpacketization and slightly higher load compare to some existing schemes with exponential subpacketization.

### v2

@@ V-A3 Comparison with the WCWL scheme in [25]
Let $K=q^{n}L$ and $t=q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}$ in [25]. We compare the WCWL scheme with the proposed scheme in the following cases:
$g_{\text{WCWL}}$ $=2\lfloor\frac{K}{K-tL+1}\rfloor=2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor.$
@@
$\frac{g_{\text{WCWL}}}{g}$ $=\frac{2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor}{2^{n}}=2^{1-n}\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor$ (17)
@@
    $F_{\text{WCWL}}=K=q^{n}L,R_{\text{WCWL}}=\frac{(K-tL)(K-tL+1)}{2K}=\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\left(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1\right)}{2q^{n}}.$
@@
    $\frac{F_{\text{WCWL}}}{F}=\frac{q^{n}L}{q^{n}L}=1,\quad\frac{R_{\text{WCWL}}}{R}=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\left(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1\right)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=2^{n}\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}{2q^{n}L}\textgreater\frac{2^{n}(q-1)^{n}}{2q^{n}}=\frac{1}{2}\cdot(2-\frac{2}{q})^{n}.$
@@
    $F_{\text{WCWL}}=\left(2\lfloor\frac{K}{K-tL+1}\rfloor+1\right)K,\ \ \ R_{\text{WCWL}}=\frac{K-tL}{2\lfloor\frac{K}{K-t+1}\rfloor+1}=\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor+1}.$
@@
    $\frac{R_{\text{WCWL}}}{R}$ $=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor+1}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}}{2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor+1}>\frac{2^{n}(q-3)^{n}L}{2q^{n}L+(q-1)^{n}L+1}>\frac{2^{n}(q-3)^{n}}{3q^{n}}=\frac{1}{3}\left(2-\frac{1}{6}\right)^{n}.$
  • Otherwise, the WCWL scheme can be obtained as follow.
    $F_{\text{WCWL}}=2\left\lfloor\frac{K}{K-tL+1}\right\rfloor K,\ \ \ R_{\text{WCWL}}=\frac{K-tL}{2\lfloor\frac{K}{K-tL+1}\rfloor}={\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor}}.$
@@
    $\frac{F_{\text{WCWL}}}{F}=\frac{2\left\lfloor\frac{K}{K-tL+1}\right\rfloor K}{q^{n}L}=2\left\lfloor\frac{K}{K-tL+1}\right\rfloor,\frac{R_{\text{WCWL}}}{R}=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L}{2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}}{2\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor}>\frac{2^{n}}{2(\frac{q}{{q-1}})^{n}}=\frac{1}{2}\cdot(2-\frac{2}{q})^{n}.$
As a result, compared to the WCWL scheme in [25], our scheme obtaining the same subpacketization or the reduction amount of our subpacketization is larger than $2$ times. The multiplicative reduction amount by our transmission is at least $\frac{1}{3}\left(2-\frac{1}{6}\right)^{n}$. When $q>6$ and $n\geq\frac{\ln 3}{\ln(2-\frac{6}{q})}$, the formula $\frac{1}{3}\left(2-\frac{1}{6}\right)^{n}\geq 1$ always holds. The coded caching gain of our scheme is $2^{1-n}\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor$ times that of the WCWL scheme. Since $K\geq 3^{n}L$, $2^{1-n}\left\lfloor\frac{q^{n}L}{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+1}\right\rfloor$ is always less than or equal to 1.
@@ V-A4 Comparison with the SR2 scheme in [28]
Let $K=q^{n}L$ and $k=q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}$ in [28], i.e., the $10$th row of Table I. We compare the SR2 scheme with the proposed scheme in the following cases:
$F_{\text{SR2}}=K=q^{n}L,\quad R_{\text{SR2}}=\frac{(K-kL)(K-kL+k)}{2K}=\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\left(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\right)}{2q^{n}}.$
@@
$\frac{F_{\text{SR2}}}{F}=\frac{q^{n}L}{q^{n}L}=1,\quad\frac{R_{\text{SR2}}}{R}=\frac{\frac{2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\left(2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}L+q^{n}-2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}\right)}{2q^{n}}}{\lfloor\frac{q-1}{2}\rfloor^{n}L}=\frac{2^{n}2^{n}\lfloor\frac{q-1}{2}\rfloor^{n}(L-1)+q^{n}}{2q^{n}L}>\frac{L-1}{2L}\left(2-\frac{6}{q}\right)^{n}.$
As a result, compared to the SR2 scheme in [28], our scheme either obtaining the same subpacketization. The multiplicative reduction amount by our transmission is at least $\frac{L-1}{2L}\left(2-\frac{6}{q}\right)^{n}$. When $q>6$ and $n\geq\frac{\ln\frac{2L}{L-1}}{\ln(2-\frac{6}{q})}$ ,the formula $\frac{L-1}{2L}\left(2-\frac{6}{q}\right)^{n}\geq 1$ always holds.
@@ V-B Numerical comparisons
Comparison with the schemes in [8, 22, 25, 23, 24]
When $K=84$ and $L=3$, we have our scheme and the schemes in [8, 22, 25, 23, 24] are shown in Fig. 3(a) and Fig. 3(b). It can be seen that our scheme can achieve a significant lower subpacketization and a slightly higher load than the schemes in [8, 23, 24, 22] with exponential subpaketization, a lower subpacketization and a same or lower load than the scheme in [25]. For instance, when $M/N=0.14$, we have our scheme with $F=84$ and $R=11.40$; the schemes in [8, 23, 22, 24] have lowest subpacketization is $F\approx 8.687\times{10}^{7}>84$ and lowest load is $R=4.10<11.40$. The WCWL scheme in [25] with $F\approx 170>84$ and $R\approx 27.98>11.40$.
(a) Memory ratio-subpacketization tradeoff for $K=84$, $L=3$
(b) Memory ratio-load tradeoff for $K=84$, $L=3$
@@ VI Conclusion
This paper proposes a novel combinational structure called as cyclic multi-access non-half-sum disjoint packing (CMA-NHSDP), which is an extension of the traditional NHSDP structure to the MACC system. The CMA-NHSDP can realize a multi-access coded caching scheme with linear subpacketization while minimizing transmission load. In terms of performance comparison, compared to some existing schemes with linear subpacketization, our scheme has a larger coded caching gain, thus achieving a lower rate. And our scheme achieves far lower subpacketization and slightly higher load compared to some existing schemes with exponential subpacketization.